# Topology-Driven Quantum Architecture Search Framework

### v1

@@ 5 Conclusion and Future Work
In this paper, we proposed the TD-QAS framework, which significantly enhances the efficiency and scalability of QAS by decoupling the search process into two key steps: topology searching and gate-type fine-tuning. This novel approach reduces the size and complexity of the search space, allowing for a more resource-efficient exploration. Our experiments across various tasks and scenarios, including both noisy and noiseless scenarios, demonstrate that TD-QAS can effectively discover high-performing PQCs with reduced quantum computational costs. The TD-QAS framework was successfully implemented with two widely used QAS algorithms, i.e., QuantumSupernet and DQAS, demonstrating its compatibility. The numerical results indicate that TD-QAS not only improves QAS capacity but also reduces the risk of convergence to suboptimal circuits.
Under future directions, we first aim to extend TD-QAS to other QAS algorithms, broadening its utility and adaptability. Then we will focus on optimizing the evaluation strategy, such as using a predictor-base evaluation method. Additionally, we plan to explore the scalability of TD-QAS for larger quantum circuits and more complex problem spaces. By refining these aspects, we aim to establish TD-QAS as a foundational approach for efficient and scalable QAS across various quantum computing applications.
@@
This work is supported by National Natural Science Foundation of China (Grant Nos. 62372048, 62371069, 62272056)

### v2

@@ 5 Conclusion and future work
In this study, we proposed the TD-QAS framework, which significantly enhances the efficiency and scalability of QAS by decoupling the search process into two key steps: topology searching and gate-type fine-tuning. This novel approach minimizes the size and complexity of the search space, allowing for a more resource-efficient exploration. This method represents a novel paradigm for QAS, which can easily be integrated with both existing and future QAS algorithms without altering their core mechanisms. Our experiments across various tasks, conducted in both noisy and noiseless scenarios, demonstrate that TD-QAS effectively discovers high-performing PQCs while reducing quantum computational costs. The TD-QAS framework was successfully implemented with two widely used QAS algorithms, i.e., QuantumSupernet and DQAS, demonstrating its compatibility. The numerical results indicate that TD-QAS improves QAS capacity while minimizing the risk of convergence to suboptimal circuits. To enhance reproducibility, the code for TD-QAS and all experiments will be made publicly available upon acceptance at: https://github.com/Sujun124/TD-QAS.
In future work, we first aim to extend TD-QAS to other QAS algorithms, thereby enhancing its utility and adaptability. Next, we will focus on optimizing the evaluation strategy, such as implementing a predictor-based evaluation method. Additionally, we plan to explore the scalability of TD-QAS for more complex tasks and scenarios. While our current study validates the core assumption and effectiveness of TD-QAS on three representative VQA tasks under both noiseless and noisy scenarios, this validation is limited to quantum machine learning tasks. Within the tasks explored, gate-type selection did not appear to play an equally or more critical role than topology. However, this assumption may not hold across all quantum computing settings. Investigating a broader range of tasks and environments remains an interesting direction for future research. By refining these aspects, we aim to establish TD-QAS as a foundational approach for efficient and scalable QAS across various quantum computing applications.
@@
This work is supported by the National Natural Science Foundation of China (Grant Nos. 62372048, 62371069, 62272056)